OpenStudy (anonymous):

which could be converted to an exponential equation? a.x=log(2) 8 b.3=log(x)8 c.3=log(2)x d.3=log(2)8

OpenStudy (blackstreet23):

are those powers?

OpenStudy (solomonzelman):

any of them could be

OpenStudy (solomonzelman):

but, "which SHOULD be converted" is a better question

OpenStudy (solomonzelman):

For example option (a). \(\color{#000000 }{ \displaystyle x=\log_28=\log_2(2^3)=3\times \log_22=3\times 1=3 }\) So the x is already given there, once you apply the logarithmic properties. Although you COULD convert to exponential form if you really like, \(\color{#000000 }{ \displaystyle x=\log_28\quad \Longleftrightarrow \quad 2^x=8 }\)

OpenStudy (anonymous):

okay thanks :)